# A convergence proof for a finite element discretization of Chorin’s projection method of the incompressible Navier-Stokes equations

### v1

@@ Abstract.
We study Chorin’s projection method combined with a finite element spatial discretization for the time-dependent incompressible Navier–Stokes equations. The scheme advances the solution in two steps: a prediction step, which computes an intermediate velocity field that is generally not divergence-free, and a projection step, which enforces incompressibility by projecting this velocity onto the divergence-free subspace. We establish convergence, up to a subsequence, of the numerical approximations generated by this scheme to a Leray–Hopf weak solution of the Navier–Stokes equations, without any additional regularity assumptions beyond square-integrable initial data. A discrete energy inequality yields a priori estimates, which we combine with a new compactness result to prove precompactness of the approximations in $L^{2}([0,T]\times\Omega)$, where $[0,T]$ is the time interval and $\Omega$ is the spatial domain. Passing to the limit as the discretization parameters vanish, we obtain a weak solution of the Navier–Stokes equations. A central difficulty is that different a priori bounds are available for the intermediate and projected velocity fields; our compactness argument carefully integrates these estimates to complete the convergence proof.
@@ 1. Introduction
Despite their long history, rigorous convergence analyses of projection methods to weak solutions of the Navier–Stokes equations, particularly in fully discrete settings and for nonsmooth data, have only recently gained traction. Classical analyses often assume high regularity of the exact solution or are limited to semi-discrete (time-discrete) schemes [16, 18, 20, 21, 1, 9, 26, 28, 30, 17]. In three spatial dimensions, where uniqueness and regularity are open problems, convergence to a Leray–Hopf weak solution (existence guaranteed by Leray [23] and Hopf [19]) is the natural benchmark.
Fully discrete convergence results have been established over the last few years for finite difference schemes on arbitrary Lipschitz domains [22, 25], and for finite volume methods on rectangular domains [13]. More recently, Eymard and Maltese [12] proved convergence of an incremental projection finite element method under the assumption of conforming finite element spaces, relying on a careful construction of a specialized interpolation operator. However, no convergence proof existed so far for the original Chorin projection method discretized via finite elements using the Darcy-type projection formulation.
Our work fills this gap by providing a straightforward and general proof of convergence for a fully discrete finite element scheme based on the original Chorin projection method with a Darcy formulation of the projection step. Our analysis covers Lipschitz domains in two and three dimensions with minimal regularity assumptions on the initial data and forcing. The main difficulty arises from the presence of two discrete velocity approximations at each time step: the intermediate velocity $\widetilde{u}$ enjoys spatial $H^{1}$-bounds, while the corrected velocity $u$ satisfies time-derivative bounds but only $L^{2}$ spatial control. Classical compactness tools, such as the Aubin–Lions–Simon lemma, cannot be applied directly because these bounds reside on different variables and no single variable carries enough regularity in both space and time. To overcome this, we prove a novel compactness result (Lemma 4.4) that combines the complementary regularity of $\widetilde{u}$ and $u$ to establish strong precompactness in $L^{2}((0,T)\times\Omega)$. This new lemma represents the core technical advance of our paper and could be of independent interest for other fractional step methods or multiphysics systems with split variables. Utilizing this tool, we show convergence (up to a subsequence) of the finite element approximations to a Leray–Hopf weak solution of the Navier–Stokes equations as both the spatial mesh size and time step tend to zero.
@@
The remainder of this paper is organized as follows. In Section 2, we introduce the necessary notation and recall the notion of Leray–Hopf weak solutions. Section 3 presents the fully discrete finite element scheme for the Chorin projection method and establishes key a priori estimates. Section 4 contains the proof of convergence to a weak solution, relying on our new compactness lemma. In Section 5, we adapt the convergence proof of the incremental projection scheme from [12] to a more general setting. Auxiliary results are collected in the Appendix.
@@ 2.1. Notation
$L^{2}_{0}(\Omega)=\{\phi\in L^{2}(\Omega);\,\int_{\Omega}\phi dx=0\},$
@@
Assume that $u_{0}\in L^{2}_{\operatorname{div}}(\Omega)$ and $f\in L^{2}(0,T;L^{2}_{\operatorname{div}}(\Omega))$. Let $u:[0,T]\times\Omega\to\mathbb{R}^{d}$ be a vector field that is divergence-free, i.e., $\operatorname{div}u=0$ for almost every $(t,x)\in[0,T]\times\Omega$ and that satisfies
@@
Moreover, $u$ satisfies the distributional version of (1.1):
(2.3) $\int_{0}^{T}({u},\partial_{t}v)dt+(u_{0},v(0,\cdot))+\int_{0}^{T}\int_{\Omega}((u\cdot\nabla)v)\cdot udxdt=-\int_{0}^{T}(f,v)dt+\mu\int_{0}^{T}(\nabla u,\nabla v)dt,$
for test functions $v\in C_{c}^{\infty}([0,T);C^{\infty}_{c,\operatorname{div}}(\Omega))$. Moreover, let $u$ satisfy the energy inequality:
@@ 3.1. Time discretization
Specifically, we let $N\in\mathbb{N}$ the number of time levels and discretize the time interval $[0,T]$ into time levels $t^{m}=m\Delta t$, $m=0,1,\dots,N$ where $\Delta t=T/N$. At each time level $t^{m}$, we seek discrete approximations $u^{m}(x)$ and $p^{m}(x)$ to $u(t^{m},x)$ and $p(t^{m},x)$. We denote the initial data $(u^{0},p^{0})=(u_{0},0)$. Then at every time level, given $(u^{m},p^{m})$, for $m\in 0,\dots,M-1$, we update $(u^{m+1},p^{m+1})$ according to the following two steps:
@@ 2. Step 2
(3.5) $\frac{1}{2}\left\lVert u^{N}\right\rVert^{2}_{L^{2}}+\frac{1}{2}\sum_{m=0}^{N-1}\left\lVert u^{m+1}-\widetilde{u}^{m+1}\right\rVert_{L^{2}}^{2}+\frac{1}{2}\sum_{m=0}^{N-1}\left\lVert\widetilde{u}^{m+1}-u^{m}\right\rVert_{L^{2}}^{2}\\ +\mu\sum_{m=0}^{N-1}\left\lVert\nabla\widetilde{u}^{m+1}\right\rVert_{L^{2}}^{2}\Delta t\leq Ce^{ct^{N}}\left(\frac{1}{2}\left\lVert u^{0}\right\rVert^{2}_{L^{2}}+\frac{\Delta t}{2}\sum_{m=0}^{N-1}\left\lVert f^{m}\right\rVert_{L^{2}}^{2}\right),$
@@ 3.2. Spatial discretization
Next we describe the spatial discretization for the problem. We let $\mathcal{T}_{h}=\{K\}$ be families of conforming quasi-uniform triangulations of $\Omega$ made of simplices with mesh size $h>0$. Here $h=\max_{K\in\mathcal{T}_{h}}\text{diam}(K)$. For simplicity, we assume that $\Omega=\Omega_{h}$, so that there is no geometric error caused by the domain approximation. We let $\mathbb{U}_{h}\subset H^{1}_{0}(\Omega)^{d}$, and $\mathbb{P}_{h}\subset L^{2}_{0}(\Omega)$ be finite dimensional subspaces scaled by a meshsize $h>0$ that we will use for the spatial approximation of the velocity, and the pressure respectively. The finite element spaces $\mathbb{U}_{h}$ and $\mathbb{P}_{h}$ need to satisfy the following two requirements: First, we require the $L^{2}$-orthgonal projection $\mathcal{P}_{\mathbb{U}_{h}}:L^{2}(\Omega)^{d}\to\mathbb{U}_{h}$ defined by
(3.6) $(u,\mathcal{P}_{\mathbb{U}_{h}}v)=(u,v),\quad\forall\,u\in\mathbb{U}_{h},$
to be stable on $H^{2}$ and $L^{r}$, $1\leq r\leq\infty$, and have good approximation properties. In particular, we need for some constant $C>0$,
@@
where $k\in\mathbb{N}$ is the order of the finite element space. Under the assumption of quasi-uniformity of the meshes, (3.7) can be achieved for finite elements based on piecewise polynomial spaces such as
@@
where $\beta^{*}$ is independent of the discretization parameter $h>0$. This is crucial for obtaining a stable approximation of the velocity-pressure pair of the incompressible Navier-Stokes equations (see e.g. [3, 4, 18] and references therein). For examples of pairs of finite element spaces $(\mathbb{U}_{h},\mathbb{P}_{h})$ that satisfy the LBB condition (3.8), see for example [14, 10]. The results that we present in the following apply to any pair of finite element spaces $(\mathbb{U}_{h},\mathbb{P}_{h})$ satisfying (3.7) and (3.8).
@@
where $\mathcal{P}_{\mathbb{U}_{h}}$ is the $L^{2}$-orthogonal projection. Since $\widetilde{u}_{h}^{0}$ may not be divergence free, we project it to a divergence free function using the following projection step: Find $(u_{h}^{0},p_{h}^{0})\in\mathbb{U}_{h}\times\mathbb{P}_{h}$ such that
@@ 1. Step 1
    (Prediction step): For any $m\geq 0$, given $u_{h}^{m}\in\mathbb{U}_{h}$, find $\widetilde{u}_{h}^{m+1}\in\mathbb{U}_{h}$, such that for all $v\in\mathbb{U}_{h}$,
    (3.11) $\left(\frac{\widetilde{u}^{m+1}_{h}-u^{m}_{h}}{\Delta t},v\right)+b(u^{m}_{h},\widetilde{u}^{m+1}_{h},v)+\mu(\nabla\widetilde{u}_{h}^{m+1},\nabla v)=(\mathcal{P}_{\mathbb{U}_{h}}f^{m},v),$
@@ 2. Step 2
Remark $3.$
It follows from using an LBB-stable pair $\mathbb{U}_{h}\times\mathbb{P}_{h}$ that the operator $B_{h}:\mathbb{U}_{h}\to\mathbb{P}_{h}$ defined by
$(B_{h}v,q)=b(v,q)=-(\operatorname{div}v,q)=(v,B_{h}^{\top}q),\quad\forall\,q\in\mathbb{P}_{h},\,v\in\mathbb{U}_{h},$
is surjective [18, 3, 14]. Here $B_{h}^{T}:\mathbb{P}_{h}\to\mathbb{U}_{h}$ is the adjoint operator of $B_{h}$. Thus we can take $q=\operatorname{div}u^{m+1}_{h}\in\mathbb{U}_{h}$ as a test function in (3.12b) to obtain
(3.13) $\left\lVert\operatorname{div}u_{h}^{m+1}\right\rVert_{L^{2}}^{2}=0,$
which implies that $\operatorname{div}u_{h}^{m+1}=0$ a.e. in $\Omega$. Hence $u_{h}^{m}\in L^{2}_{\operatorname{div}}(\Omega)$, $m=1,\dots,N$.
@@ 3.3. Solvability of the scheme
Lemma 3.4.
For every $\Delta t,h>0$ and every $m=0,1,\dots$, given $(u^{m}_{h},p^{m}_{h})\in\mathbb{U}_{h}\times\mathbb{P}_{h}$, then there exists a unique $(u^{m+1}_{h},p^{m+1}_{h})\in\mathbb{U}_{h}\times\mathbb{P}_{h}$ solving (3.11)–(3.12).
@@
We start by showing that if $(u_{h}^{m},p^{m}_{h})\in\mathbb{U}_{h}\times\mathbb{P}_{h}$, then there exists a unique $\widetilde{u}^{m+1}_{h}\in\mathbb{U}_{h}$ satisfying (3.11), i.e., the first step of the algorithm is well-posed. To do so, we write (3.11) as a linear variational problem: Find $\widetilde{u}\in\mathbb{U}_{h}$ such that for all $v\in\mathbb{U}_{h}$,
@@
(3.15) $a^{m}(\widetilde{u},v)=\left(\widetilde{u},v\right)+\Delta tb(u^{m}_{h},\widetilde{u},v)+\Delta t\mu(\nabla\widetilde{u},\nabla v)$
@@
$a^{m}(\widetilde{u},\widetilde{u})=\left\lVert\widetilde{u}\right\rVert_{L^{2}}^{2}+\Delta tb(u^{m}_{h},\widetilde{u},\widetilde{u})+\Delta t\mu(\nabla\widetilde{u},\nabla\widetilde{u})=\left\lVert\widetilde{u}\right\rVert_{L^{2}}^{2}+\Delta t\mu\left\lVert\nabla\widetilde{u}\right\rVert_{L^{2}}^{2}\geq c\left\lVert\widetilde{u}\right\rVert_{H^{1}_{0}}^{2},$
@@
$\left|a^{m}(\widetilde{u},v)\right|\leq$ $\left\lVert\widetilde{u}\right\rVert_{L^{2}}\left\lVert v\right\rVert_{L^{2}}+\Delta t\left\lVert u^{m}_{h}\right\rVert_{L^{4}}\left\lVert\widetilde{u}\right\rVert_{H^{1}_{0}}\left\lVert v\right\rVert_{L^{4}}+\Delta t\mu\left\lVert\nabla\widetilde{u}\right\rVert_{L^{2}}\left\lVert\nabla v\right\rVert_{L^{2}}$
$\leq$ $\left\lVert\widetilde{u}\right\rVert_{L^{2}}\left\lVert v\right\rVert_{L^{2}}+\Delta t(\left\lVert u^{m}_{h}\right\rVert_{L^{4}}+\mu)\left\lVert\widetilde{u}\right\rVert_{H^{1}_{0}}\left\lVert v\right\rVert_{H^{1}_{0}}\leq C\left\lVert\widetilde{u}\right\rVert_{H^{1}_{0}}\left\lVert v\right\rVert_{H^{1}_{0}},$
where we used the Sobolev embedding theorem and the Poincaré inequality. Here we also used that $u^{m}_{h}\in H^{1}_{0}(\Omega)$ by assumption (and induction). Hence $a^{m}$ is bounded on $\mathbb{U}_{h}$. To show boundedness of $\mathcal{F}^{m}$ (for every $m$), we estimate:
@@
The solvability of Step 2 follows from the choice of the LBB-stable pair $(\mathbb{U}_{h},\mathbb{P}_{h})$: Corollary 3.1 in [16] implies the orthogonal decomposition $\mathbb{U}_{h}=\mathbb{H}_{h}\oplus B_{h}^{\top}(\mathbb{P}_{h}^{1})$, where $\mathbb{H}_{h}=\ker(B_{h})$ and where $\mathbb{P}_{h}^{1}$ is the subset of $q\in\mathbb{P}_{h}$ for which $\left\lVert q\right\rVert_{\mathbb{P}_{h}^{1}}=\left\lVert B_{h}^{\top}q\right\rVert_{L^{2}(\Omega)}$ is bounded. (Here $\left\lVert\cdot\right\rVert_{\mathbb{P}_{h}^{1}}$ is induced by the inner product $(p,q)_{\mathbb{P}_{h}^{1}}=(B_{h}^{\top}p,B_{h}^{1}q)$). Thus any $v\in\mathbb{U}_{h}$ can be decomposed into $v=w+B_{h}^{\top}p$, where $p\in\mathbb{P}_{h}^{1}$ and $w\in\mathbb{U}_{h}$ satisfies $B_{h}w=0,$ and $w$ and $B_{h}^{\top}p$ are orthogonal. This decomposition is unique. To see this, assume by contradiction that there are $w_{1},w_{2}\in\mathbb{H}_{h}$ and $p_{1},p_{2}\in\mathbb{P}_{h}^{1}$ such that $v=w_{1}+B_{h}^{\top}p_{1}=w_{2}+B_{h}^{\top}p_{2}$. We subtract the two equations from each other to obtain
@@
$\begin{split}\widetilde{u}^{m+1}_{h}&=u^{m+1}_{h}+B_{h}^{\top}p_{h}^{m+1},\\ B_{h}u^{m+1}_{h}&=0,\end{split}$
i.e., $u^{m+1}_{h}\in\ker(B_{h})=\mathbb{H}_{h}$ and $B_{h}^{\top}p_{h}^{m+1}\in B_{h}^{\top}(\mathbb{P}_{h}^{1})$ with $\widetilde{u}^{m+1}_{h}=u^{m+1}_{h}+B_{h}^{\top}p_{h}^{m+1}$ and we already know this decomposition is unique (see also [18]). ∎

### v2

@@ Abstract.
We study Chorin’s projection method combined with a finite element spatial discretization for the time-dependent incompressible Navier–Stokes equations. The scheme advances the solution in two steps: a prediction step, which computes an intermediate velocity field that is generally not divergence-free, and a projection step, which enforces (approximate) incompressibility by projecting this velocity onto the (approximately) divergence-free subspace. We establish convergence, up to a subsequence, of the numerical approximations generated by this scheme to a Leray–Hopf weak solution of the Navier–Stokes equations, without any additional regularity assumptions beyond square-integrable initial data. A discrete energy inequality yields a priori estimates, which we combine with a new compactness result to prove precompactness of the approximations in $L^{2}([0,T]\times\Omega)$, where $[0,T]$ is the time interval and $\Omega$ is the spatial domain. Passing to the limit as the discretization parameters vanish, we obtain a weak solution of the Navier–Stokes equations. A central difficulty is that different a priori bounds are available for the intermediate and projected velocity fields; our compactness argument carefully integrates these estimates to complete the convergence proof.
@@ 1. Introduction
Despite their long history, rigorous convergence analyses of projection methods to weak solutions of the Navier–Stokes equations, particularly in fully discrete settings and for nonsmooth data, have only recently gained traction. Classical analyses often assume high regularity of the exact solution or are limited to semi-discrete (time-discrete) schemes [16, 18, 20, 21, 1, 9, 25, 27, 29, 17]. In three spatial dimensions, where uniqueness and regularity are open problems, convergence to a Leray–Hopf weak solution (existence is guaranteed by Leray [23] and Hopf [19]) is the natural benchmark.
Fully discrete convergence results have been established over the last few years for finite difference schemes on arbitrary Lipschitz domains [22, 24], and for finite volume methods on rectangular domains [13]. More recently, Eymard and Maltese [12] proved convergence of an incremental projection finite element method with Poisson formulation of the projection step under the assumption of conforming finite element spaces, relying on a careful construction of a specialized interpolation operator. However, no convergence proof existed so far for the original Chorin projection method discretized via finite elements using the Darcy-type projection formulation.
Our work fills this gap by providing a proof of convergence for a fully discrete finite element scheme based on the original Chorin projection method with a Darcy formulation of the projection step. Our analysis covers Lipschitz domains in two and three dimensions with minimal regularity assumptions on the initial data and forcing. The main difficulty arises from the lack of sufficient bounds on the pressure approximation and the intermediate velocity $\widetilde{u}$ not being divergence free. So we can only obtain a bound on the time derivative of the approximation of $\widetilde{u}$ in the dual space of $H^{1}_{\operatorname{div}}\cap H^{k}(\Omega)$ (where $k\in\mathbb{N}$ is sufficiently large and $H^{1}_{\operatorname{div}}(\Omega)$ denotes the space of almost everywhere divergence free $H^{1}_{0}(\Omega)$ functions). Therefore, classical compactness results such as the Aubin-Lions-Simon lemma cannot be applied directly. To overcome this, we take a new approach to precompactness that carefully combines the a priori estimates on the approximation of the intermediate velocity $\widetilde{u}$ with the approximately divergence free property of the approximation of the corrected velocity $u$. This way, we achieve strong precompactness in $L^{2}([0,T]\times\Omega)$ for the approximations of $\widetilde{u}$ and $u$. This new approach represents the core technical advance of our paper and could be of independent interest for other fractional step methods or multiphysics systems with split variables. Utilizing this tool, we show convergence (up to a subsequence) of the finite element approximations to a Leray–Hopf weak solution of the Navier–Stokes equations as both the spatial mesh size and time step tend to zero.
@@
The remainder of this paper is organized as follows. In Section 2, we introduce the necessary notation and recall the notion of Leray–Hopf weak solutions. Section 3 presents the fully discrete finite element scheme for the Chorin projection method and establishes key a priori estimates. Section 4 contains the proof of convergence to a weak solution. In Section 5, we adapt the convergence proof of the incremental projection scheme from [12] to a more general setting. Auxiliary results are collected in the Appendix.
@@ 2.1. Notation
$L^{2}_{0}(\Omega)=\{\phi\in L^{2}(\Omega);\,\int_{\Omega}\phi dx=0\}$
@@
Assume that $u_{0}\in L^{2}_{\operatorname{div}}(\Omega)$ and $f\in L^{2}(0,T;L^{2}_{\operatorname{div}}(\Omega))$. Let $u:[0,T]\times\Omega\to\mathbb{R}^{d}$ be a vector field that is divergence-free, i.e., $\operatorname{div}u=0$ for almost every $(t,x)\in[0,T]\times\Omega$, and that satisfies
@@
Assume that $u$ satisfies the distributional version of (1.1):
(2.3) $\int_{0}^{T}({u},\partial_{t}v)dt+\int_{0}^{T}\int_{\Omega}((u\cdot\nabla)v)\cdot udxdt=-\int_{0}^{T}(f,v)dt+\mu\int_{0}^{T}(\nabla u,\nabla v)dt,$
for test functions $v\in C_{c}^{\infty}((0,T);C^{\infty}_{c,\operatorname{div}}(\Omega))$. Moreover, let $u$ satisfy the energy inequality:
@@ 3.1. Time discretization
Specifically, we let $N\in\mathbb{N}$ the number of time levels and discretize the time interval $[0,T]$ into time levels $t^{m}=m\Delta t$, $m=0,1,\dots,N$ where $\Delta t=T/N$. At each time level $t^{m}$, we seek discrete approximations $u^{m}(x)$ and $p^{m}(x)$ to $u(t^{m},x)$ and $p(t^{m},x)$. We denote the initial data $(u^{0},p^{0})=(u_{0},0)$. Then at every time level $m=0,\dots,M-1$, given $(u^{m},p^{m})$, we update $(u^{m+1},p^{m+1})$ according to the following two steps:
@@ 2. Step 2
(3.5) $\frac{1}{2}\left\lVert u^{M}\right\rVert^{2}_{L^{2}}+\frac{1}{2}\sum_{m=0}^{M-1}\left\lVert u^{m+1}-\widetilde{u}^{m+1}\right\rVert_{L^{2}}^{2}+\frac{1}{2}\sum_{m=0}^{M-1}\left\lVert\widetilde{u}^{m+1}-u^{m}\right\rVert_{L^{2}}^{2}\\ +\mu\sum_{m=0}^{M-1}\left\lVert\nabla\widetilde{u}^{m+1}\right\rVert_{L^{2}}^{2}\Delta t\leq Ce^{ct^{M}}\left(\frac{1}{2}\left\lVert u^{0}\right\rVert^{2}_{L^{2}}+\frac{\Delta t}{2}\sum_{m=0}^{M-1}\left\lVert f^{m}\right\rVert_{L^{2}}^{2}\right),$
@@ 3.2. Spatial discretization
Next we describe the spatial discretization for the problem. We let $\mathcal{T}_{h}=\{K\}$ be families of conforming quasi-uniform triangulations of $\Omega$ made of simplices with mesh size $h>0$. Here $h=\max_{K\in\mathcal{T}_{h}}\text{diam}(K)$. For simplicity, we assume that $\Omega=\Omega_{h}$, so that there is no geometric error caused by the domain approximation. We let $\mathbb{U}_{h}\subset H^{1}_{0}(\Omega)^{d}$, and $\mathbb{P}_{h}\subset L^{2}_{0}(\Omega)\cap H^{1}(\Omega)$ be finite dimensional subspaces scaled by a meshsize $h>0$ that we will use for the spatial approximation of the velocity, and the pressure respectively. The finite element spaces $\mathbb{U}_{h}$ and $\mathbb{P}_{h}$ need to satisfy the following two requirements: First, we require the $L^{2}$-orthgonal projections $\mathcal{P}_{\mathbb{U}_{h}}:L^{2}(\Omega)^{d}\to\mathbb{U}_{h}$, $\mathcal{P}_{\mathbb{P}_{h}}:L^{2}(\Omega)\to\mathbb{P}_{h}$ defined by
(3.6) $(u,\mathcal{P}_{\mathbb{U}_{h}}v)=(u,v),\quad\forall\,u\in\mathbb{U}_{h},\quad(p,\mathcal{P}_{\mathbb{U}_{h}}q)=(p,q),\quad\forall\,p\in\mathbb{P}_{h},$
to be stable on $H^{1}$ and $L^{r}$, $1\leq r\leq\infty$, and have good approximation properties. In particular, we need for some constant $C>0$, independent of the mesh size $h>0$,
@@
where $k\in\mathbb{N}$ is the order of the finite element space, and we require the same properties for $\mathcal{P}_{\mathbb{P}_{h}}$ for a possibly different $k$. Under the assumption of quasi-uniformity of the meshes, (3.7) can be achieved for finite elements based on piecewise polynomial spaces such as
@@
where $\beta^{*}$ is independent of the discretization parameter $h>0$. This is crucial for obtaining a stable approximation of the velocity-pressure pair of the Stokes equations (see e.g. [4, 5, 18] and references therein). For examples of pairs of finite element spaces $(\mathbb{U}_{h},\mathbb{P}_{h})$ that satisfy the LBB condition (3.8), see for example [14, 10]. The results that we present in the following apply to any pair of finite element spaces $(\mathbb{U}_{h},\mathbb{P}_{h})\subset H^{1}_{0}(\Omega)^{d}\times L^{2}_{0}(\Omega)\cap H^{1}(\Omega)$ satisfying (3.7) and (3.8). One popular option are the Taylor-Hood finite elements, but also the MINI element by Arnold, Brezzi and Fortin [2] satisfies the requirements.
@@
where $\mathcal{P}_{\mathbb{U}_{h}}$ is the $L^{2}$-orthogonal projection. Since $\widetilde{u}_{h}^{0}$ may not be divergence free, we project it to an approximately divergence free function using the following projection step: Find $(u_{h}^{0},p_{h}^{0})\in\mathbb{U}_{h}\times\mathbb{P}_{h}$ such that
@@ 1. Step 1
    (Prediction step): For any $m\geq 0$, given $u_{h}^{m},\widetilde{u}_{h}^{m}\in\mathbb{U}_{h}$, find $\widetilde{u}_{h}^{m+1}\in\mathbb{U}_{h}$, such that for all $v\in\mathbb{U}_{h}$,
    (3.11) $\left(\frac{\widetilde{u}^{m+1}_{h}-u^{m}_{h}}{\Delta t},v\right)+b(\widetilde{u}^{m}_{h},\widetilde{u}^{m+1}_{h},v)+\mu(\nabla\widetilde{u}_{h}^{m+1},\nabla v)=(\mathcal{P}_{\mathbb{U}_{h}}f^{m},v),$
@@ 3.3. Solvability of the scheme
Lemma $3.$
For every $\Delta t,h>0$ and every $m=0,1,\dots$, given $(\widetilde{u}_{h}^{m},u^{m}_{h},p^{m}_{h})\in\mathbb{U}_{h}\times\mathbb{U}_{h}\times\mathbb{P}_{h}$, then there exists a unique $(\widetilde{u}_{h}^{m+1},u^{m+1}_{h},p^{m+1}_{h})\in\mathbb{U}_{h}\times\mathbb{U}_{h}\times\mathbb{P}_{h}$ solving (3.11)–(3.12).
@@
We start by showing that if $(u_{h}^{m},\widetilde{u}_{h}^{m},p^{m}_{h})\in\mathbb{U}_{h}\times\mathbb{U}_{h}\times\mathbb{P}_{h}$, then there exists a unique $\widetilde{u}^{m+1}_{h}\in\mathbb{U}_{h}$ satisfying (3.11), i.e., the first step of the algorithm is well-posed. To do so, we write (3.11) as a linear variational problem: Find $\widetilde{u}\in\mathbb{U}_{h}$ such that for all $v\in\mathbb{U}_{h}$,
@@
(3.14) $a^{m}(\widetilde{u},v)=\left(\widetilde{u},v\right)+\Delta tb(\widetilde{u}^{m}_{h},\widetilde{u},v)+\Delta t\mu(\nabla\widetilde{u},\nabla v)$
@@
$a^{m}(\widetilde{u},\widetilde{u})=\left\lVert\widetilde{u}\right\rVert_{L^{2}}^{2}+\Delta tb(\widetilde{u}^{m}_{h},\widetilde{u},\widetilde{u})+\Delta t\mu(\nabla\widetilde{u},\nabla\widetilde{u})=\left\lVert\widetilde{u}\right\rVert_{L^{2}}^{2}+\Delta t\mu\left\lVert\nabla\widetilde{u}\right\rVert_{L^{2}}^{2}\geq c\left\lVert\widetilde{u}\right\rVert_{H^{1}_{0}}^{2},$
@@
$\left|a^{m}(\widetilde{u},v)\right|\leq$ $\left\lVert\widetilde{u}\right\rVert_{L^{2}}\left\lVert v\right\rVert_{L^{2}}+\Delta t\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{4}}\left\lVert\widetilde{u}\right\rVert_{H^{1}_{0}}\left\lVert v\right\rVert_{L^{4}}+\Delta t\mu\left\lVert\nabla\widetilde{u}\right\rVert_{L^{2}}\left\lVert\nabla v\right\rVert_{L^{2}}$
$\leq$ $\left\lVert\widetilde{u}\right\rVert_{L^{2}}\left\lVert v\right\rVert_{L^{2}}+\Delta t(\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{4}}+\mu)\left\lVert\widetilde{u}\right\rVert_{H^{1}_{0}}\left\lVert v\right\rVert_{H^{1}_{0}}\leq C\left\lVert\widetilde{u}\right\rVert_{H^{1}_{0}}\left\lVert v\right\rVert_{H^{1}_{0}},$
where we used the Sobolev embedding theorem and the Poincaré inequality. Here we also used that $\widetilde{u}^{m}_{h}\in H^{1}_{0}(\Omega)$ by assumption (and induction). Hence $a^{m}$ is bounded on $\mathbb{U}_{h}$. To show boundedness of $\mathcal{F}^{m}$ (for every $m$), we estimate:
@@
The solvability of Step 2 follows from the choice of the LBB-stable pair $(\mathbb{U}_{h},\mathbb{P}_{h})$: Corollary 3.1 in [16] implies the orthogonal decomposition $\mathbb{U}_{h}=\mathbb{H}_{h}\oplus B_{h}^{\top}(\mathbb{P}_{h}^{1})$, where $B_{h}:\mathbb{U}_{h}\to\mathbb{P}_{h}$ is the operator defined by
$(B_{h}v,q):=-(\operatorname{div}v,q)=(v,B_{h}^{\top}q),\quad\forall\,q\in\mathbb{P}_{h},\,v\in\mathbb{U}_{h},$
and where $\mathbb{H}_{h}=\ker(B_{h})$ and $\mathbb{P}_{h}^{1}$ is the subset of $q\in\mathbb{P}_{h}$ for which $\left\lVert q\right\rVert_{\mathbb{P}_{h}^{1}}=\left\lVert B_{h}^{\top}q\right\rVert_{L^{2}(\Omega)}$ is bounded. (Here $\left\lVert\cdot\right\rVert_{\mathbb{P}_{h}^{1}}$ is induced by the inner product $(p,q)_{\mathbb{P}_{h}^{1}}=(B_{h}^{\top}p,B_{h}^{\top}q)$). Thus any $v\in\mathbb{U}_{h}$ can be decomposed into $v=w+B_{h}^{\top}p$, where $p\in\mathbb{P}_{h}^{1}$ and $w\in\mathbb{U}_{h}$ satisfies $B_{h}w=0,$ and $w$ and $B_{h}^{\top}p$ are orthogonal. This decomposition is unique. To see this, assume by contradiction that there are $w_{1},w_{2}\in\mathbb{H}_{h}$ and $p_{1},p_{2}\in\mathbb{P}_{h}^{1}$ such that $v=w_{1}+B_{h}^{\top}p_{1}=w_{2}+B_{h}^{\top}p_{2}$. We subtract the two equations from each other to obtain
@@
$\begin{split}\widetilde{u}^{m+1}_{h}&=u^{m+1}_{h}+\Delta tB_{h}^{\top}p_{h}^{m+1},\\ B_{h}u^{m+1}_{h}&=0,\end{split}$
i.e., $u^{m+1}_{h}\in\ker(B_{h})=\mathbb{H}_{h}$ and $B_{h}^{\top}p_{h}^{m+1}\in B_{h}^{\top}(\mathbb{P}_{h}^{1})$ with $\widetilde{u}^{m+1}_{h}=u^{m+1}_{h}+\Delta tB_{h}^{\top}p_{h}^{m+1}$ and we already know this decomposition is unique (see also [18]). ∎